OpenStudy (anonymous):

What is the slope of the line that passes through (1, 4) and (1, –3)? I got 7/0. That would mean it's 0? Or undefined?

OpenStudy (anonymous):

that is right. but we canot divide by zero.

OpenStudy (anonymous):

it is -7/0

OpenStudy (anonymous):

undefined

OpenStudy (anonymous):

it is also -7/0

OpenStudy (anonymous):

So wait, it's undefined?

OpenStudy (anonymous):

no specific answer, like that. cannot solve, or syntax error. or math error.

OpenStudy (anonymous):

that is correct. the line is undefined

OpenStudy (amistre64):

the slope is defined, its just that the number produced, if viewed solely as a number is undefined. The slope is a vertical slope

OpenStudy (anonymous):

hey the answer is here the slope of the line is -7/0 so it is infinite so it is a line parallel to y-axis. because slope of a line = tan ( angle made by lione with x) for a line parallel to y the slope = tan 90 = infinite

OpenStudy (anonymous):

abhinav is right.

OpenStudy (anonymous):

lorda mercy here we go again. it is not "infinite" it is just not a number

OpenStudy (anonymous):

but he is right that it is paralel to the y-axis.

OpenStudy (anonymous):

the slope measures something. it tells you how many units up or down in the y direction the line travels as it travels one unit in the x - direction

OpenStudy (anonymous):

hey anything anything divided by zero is considered as infinite dont u know that?

OpenStudy (anonymous):

since a vertical line does not travel one unit in the x - direction, the question makes no sense and so there is no slope. like asking what is the gas mileage for my bicycle.

OpenStudy (anonymous):

really? yes i don't know it because it is not true

OpenStudy (anonymous):

even as a limit statement it is not true

OpenStudy (anonymous):

if some math teacher told you that please tell them that they are sadly mistaken

OpenStudy (anonymous):

hey dont give silly explanations what do you think i the slope of any line parallel to y axis dont u know that slope = tan (theta

OpenStudy (anonymous):

well if the slope is a number then it would be \[\tan(\theta)\] yes, but if \[\theta=\frac{\pi}{2}\] then \[\tan(\frac{\pi}{2})\] is not a number since \[\frac{\pi}{2}\] is not in the domain of tangent

OpenStudy (anonymous):

the slope has a specific meaning, as it measures something. if there is nothing to measure then it does not have a meaning

OpenStudy (anonymous):

ok so u say there is no slope for any line parallel to y axis?

OpenStudy (anonymous):

and if you were told that \[\tan(\frac{\pi}{2})=\infty\] you were told incorrectly. even as a limit this is not true

OpenStudy (anonymous):

as fas as i know the slope is infinite if it is parallel to y-axis and 0 if it is parallel to x-axis.

OpenStudy (anonymous):

far*

OpenStudy (jamesj):

Listen to satellite73. He is exactly right. The limit \[\lim_{x \rightarrow \pi/2} \tan x \] does not exist, so it makes absolutely no sense to assign a value to tan(pi/2)

OpenStudy (anonymous):

by which i mean \[\lim_{x\rightarrow \frac{\pi}{2}^+}\tan(x)=-\infty\] whereas \[\lim_{x\rightarrow \frac{\pi}{2}^-}\tan(x)=\infty\]

OpenStudy (anonymous):

ok if u dont agree with me then u can go with ur own explanation idont want to fight with u . u are not even thinking about my explanation

OpenStudy (anonymous):

i wonder where this idea comes from that the symbol \[\frac{a}{0}\]is somehow infinity. i assure you that it is not

OpenStudy (amistre64):

abhinav, your explanation is not sound; you say that the slope is infinite, which means that it moves an infinite amount when x moves, but x never moves

OpenStudy (anonymous):

satellite in some aspects it is undefined and in some aspects it is infinite.

OpenStudy (amistre64):

i have a cliff that is 3 feet high, what is its slope? infinity? no

OpenStudy (jamesj):

@abh ... Let us suppose for a moment that a/0 = inf Then what is 2a/0 - a/0 ? Is it inf - inf = 0? or is it 2inf - inf = inf? or is it inf - inf = inf? You just can't make this definition consistent.

OpenStudy (anonymous):

i do believe some math teacher may have told you this, and therefore it is in no way a fault or failing on your part. just goes to show that you should not believe everything a math teacher says.

OpenStudy (amistre64):

i vertical line has no slope that can be defined numerically; its simply a vertical slope

OpenStudy (anonymous):

of course inf -inf = inft because you will goin to subtract infintly.

OpenStudy (anonymous):

and 2inf is still inf :))

OpenStudy (anonymous):

actually the question says That would mean it's 0? Or undefined? so at least this math teacher knows something undefined is the correct answer

OpenStudy (anonymous):

if you believe \[\frac{1}{0}=\text{something}\] the presumably i can solve \[1=0\times \text{something}\] a miracle if ever there was one

OpenStudy (anonymous):

ok if u think the slope if undefined then u can go to the international maths research centre

OpenStudy (anonymous):

ok hope this will provide a correct answer since the slope of a line parallel to y axis is undefined, and 7/0 is also undefined its is a line parallel to y axis. i feel may be u are right. i am sorry for my arguement against u